# DualSG: A Dual-Stream Explicit Semantic-Guided Multivariate Time Series Forecasting Framework

### v1

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755458isbn: 979-8-4007-2035-2/2025/10ccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Many LLM-based forecasting methods can be broadly categorized into two types. LLM-only Forecasting Paradigm (Gruver et al., 2023; Tang et al., 2025b; Liu et al., 2024f; Wang et al., 2025) treat time series as text by converting them into natural language, and perform forecasting purely via LLMs. This design often leads to numerical imprecision, as continuous values must be discretized or approximated in text. It also leads to a task mismatch, as LLMs, which are inherently designed for natural language processing, are being misapplied to model time series data. Due to the limited output token budget of LLMs, this often restricts the feasible forecasting horizon. LLM-align Forecasting Paradigm (Jin et al., 2024; Liu et al., 2025; Sun et al., 2024; Pan et al., 2024) align textual and time series modalities at the token level, as illustrated in Figure 1(b). This approach frequently suffers from alignment difficulty, which distorts key time series properties, ultimately compromising the integrity of essential forecasting signals.
@@
Both paradigms thus exhibit significant shortcomings in numerical precision and alignment difficulty, raising fundamental doubts about the suitability of LLMs as forecasters for time series.
@@ 3.1. Preliminaries
[…]
@@ 3.2. Numerical Forecasting Stream
To address this, we propose Multi-scale Adaptive Patching (MAP), a non-uniform segmentation method that creates multi-scale temporal tokens.
To our knowledge, no prior patch-based time series methods (Chen et al., 2024; Wang et al., 2024a; Kudrat et al., 2025) explicitly consider temporal decay effects in patch design.
@@
[…]
@@ 3.3. Textual Reasoning Stream
[…]
@@
[…]
@@ 3.4.1. Channel Fusion via SemFuse
[…]
@@
Table 2. Computational complexity comparison among different channel fusion strategies. $N$ denotes the number of channels, and $K$ denotes the number of selected channels for CC methods.
@@
CC $O(NK)$ Compute attention weights within clusters ($K<N$) and weighted sum
@@ 4.3.2. Comparison of Prompt Strategies
Table 5. Average MAE and relative improvement (worse than DualSG) on decomposed components. DualSG achieves the most significant gain on trend modeling. Here, DualSG’s lookback windows $L=96$, so it doesn’t perform optimally on all components. The Weather dataset is used for this experiment.
@@ 4.3.3. Limitations of Fixed-length Patching and Different LLMs
[…]

### v3

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755458isbn: 979-8-4007-2035-2/2025/10ccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Many LLM-based forecasting methods can be broadly categorized into two types. LLM-only Forecasting Paradigm (Gruver et al., 2023; Tang et al., 2025b; Liu et al., 2024f; Wang et al., 2025) treat time series as text by converting them into natural language, and perform forecasting purely via LLMs. This design often leads to numerical imprecision, as continuous values must be discretized or approximated in text. It also leads to a task mismatch, as LLMs, which are inherently designed for natural language processing, are being misapplied to model time series data. Due to the limited output token budget of LLMs, this often restricts the feasible forecasting horizon. LLM-align Forecasting Paradigm (Jin et al., 2024; Liu et al., 2025; Sun et al., 2024; Pan et al., 2024) align textual and time series modalities at the token level, as illustrated in Figure 1(b). This approach frequently suffers from alignment difficulty, which distorts key time series properties, ultimately compromising the integrity of essential forecasting signals. Both paradigms thus exhibit significant shortcomings in numerical precision and alignment difficulty, raising fundamental doubts about the suitability of LLMs as forecasters for time series.
@@ 3.1. Preliminaries
Multivariate Time Series Forecasting (MTSF) aims to predict the next $F$ future time steps, given a historical multivariate time series. Formally, let the input sequence be denoted as $\mathbf{X}=\langle\mathbf{x}_{:,1},\ldots,\mathbf{x}_{:,T}$$\rangle\in\mathbb{R}^{N\times T}$, where $N$ is the number of variables (channels) and $T$ is the length of the historical window. The forecasting objective is to estimate the future sequence $\mathbf{Y}=\langle\mathbf{x}_{:,T+1},\ldots,\mathbf{x}_{:,T+F}\rangle\in\mathbb{R}^{N\times F}$.
@@ 3.2. Numerical Forecasting Stream
To address this, we propose Multi-scale Adaptive Patching (MAP), a non-uniform segmentation method that creates multi-scale temporal tokens. To our knowledge, no prior patch-based time series methods (Chen et al., 2024; Wang et al., 2024a; Kudrat et al., 2025) explicitly consider temporal decay effects in patch design.
@@
The $\text{SoftTopk}(\cdot)$ ensures differentiability, and total selection is constrained to match uniform segmentation for efficiency.
@@ 3.3. Textual Reasoning Stream
(10) $H_{\text{enc}}^{(k)}$ $=\text{TransformerBlock}(H_{\text{enc}}^{(k-1)}),\quad k=1,\dots,K$
@@
(15) $H_{\text{dec}}^{t}$ $=\text{TransformerDecoder}(E_{\text{dec}}^{t},H_{\text{align}})$
@@ 3.4.1. Channel Fusion via SemFuse
Stage 2: Sparse Top-K Selection. To promote sparsity and reduce redundancy, we construct a binary mask $\mathcal{M}_{n,i}\in\{0,1\}$ by selecting the top-$K$ most relevant channels:
@@
Table 2. Computational complexity comparison among different channel fusion strategies. NN denotes the number of channels, and $K$ denotes the number of selected channels for $C$ methods.
@@
$C$ $O(NK)$ Compute attention weights within clusters ($K<N$) and weighted sum
@@ 4.3.2. Comparison of Prompt Strategies
Table 5. Average MAE and relative improvement (% worse than DualSG) on decomposed components. DualSG achieves the most significant gain on trend modeling. Here, DualSG’s lookback windows $L=96$, so it doesn’t perform optimally on all components. The Weather dataset is used for this experiment.
@@ 4.3.3. Limitations of Fixed-length Patching and Different LLMs
Table 7. Full results for long-term forecasting consider prediction horizons $H$ within $\{96,192,336,512\}$. The lookback windows search from $\{96,192,336,512,720\}$. The term ‘Avg.” represents the average results across the four prediction lengths. The best and second best outcomes are highlighted in best and second, respectively. The notation $1^{\text{st}}$ Count” denotes the frequency of each method achieving the top results.